A school going girlis running on a horizontal road with a velocity of 4m/s. At what angle she should hold her umbrella in order to protect herself from the rains. If the rain is falling vertically with a velocity 8m/s?

Grade:11

1 Answers

Niraj Bishnoi
6 years ago
Vr2=16+64=80, or Vr=8.94
tanx=4/8.94=1/2.235=0.4474=24.10 degree
WEll you just bring tan on the other side, giving the value of x.